# Predictive Coding Networks and Inference Learning: Tutorial and Survey

### v1

@@ Abstract.
Recent years have witnessed a growing call for renewed emphasis on neuroscience-inspired approaches in artificial intelligence research, under the banner of NeuroAI. This is exemplified by recent attention gained by predictive coding networks (PCNs) within machine learning (ML). PCNs are based on the neuroscientific framework of predictive coding (PC), which views the brain as a hierarchical Bayesian inference model that minimizes prediction errors from feedback connections. PCNs trained with inference learning (IL) have potential advantages to traditional feedforward neural networks (FNNs) trained with backpropagation. While historically more computationally intensive, recent improvements in IL have shown that it can be more efficient than backpropagation with sufficient parallelization, making PCNs promising alternatives for large-scale applications and neuromorphic hardware. Moreover, PCNs can be mathematically considered as a superset of traditional FNNs, which substantially extends the range of possible architectures for both supervised and unsupervised learning. In this work, we provide a comprehensive review as well as a formal specification of PCNs, in particular placing them in the context of modern ML methods, and positioning PC as a versatile and promising framework worthy of further study by the ML community.
@@ 1. Introduction
Neuroscience-inspired approaches to Machine Learning (ML) have a long history within artificial intelligence research (rosenblatt_perceptron_1958, ; hassabis_neuroscience-inspired_2017, ; macpherson_natural_2021, ). Despite the remarkable empirical advances in ML capabilities in recent years, biological learning is still superior in many ways, such as flexibility and energy efficiency (macpherson_natural_2021, ). As such, recent years have seen a growing call for renewed emphasis on neuroscience-inspired approaches in AI research, known as NeuroAI (zador_catalyzing_2023, ; noauthor_new_2024, ). This is exemplified by the rising popularity of the predictive coding (PC, also known as predictive processing) framework in computational neuroscience (rao_predictive_1999, ; friston_free-energy_2010, ; ficco_disentangling_2021, ) and its recent entrance into the field of machine learning (whittington_approximation_2017, ; millidge_predictive_2022-1, ). PC represents a Bayesian perspective on how the brain processes information, emphasizing the role of probabilistic models and minimization of prediction errors in perception and learning (rao_predictive_1999, ; friston_hierarchical_2008, ; friston_free-energy_2010, ; ficco_disentangling_2021, ). At its core, it suggests that the brain continually generates predictions about sensory input and updates these predictions based on incoming sensory data. By removing the predictable components, this reduces redundancy in the information processing pipeline (huang_predictive_2011, ).
@@
Central to PC is the concept of hierarchical prediction: the lowest level of the hierarchy represents sensory data, and each higher level attempts to predict neural activity in the layer below (rao_predictive_1999, ). Prediction errors, arising from discrepancies between actual and predicted activity, are propagated upward through the hierarchy, while the predictions from higher levels are propagated downwards via feedback connections; see fig. 1. In the influential work by Rao & Ballard (1999), the authors showed that such a predictive coding network (PCN) can learn statistical regularities in the input and explain several neural responses in the visual cortex (shain_fmri_2020, ; okada_neural_2018, ; ficco_disentangling_2021, ; wang_predictive_2023, ; caucheteux_evidence_2023, ). Since then, the PC framework has found wide adoption across neuroscience, psychology, and philosophy (schrimpf_neural_2021, ; clark_surfing_2015, ). It is often seen as part of ‘active inference’, a broader umbrella term which aims to provide an integrated perspective on brain, cognition, and behavior (parr_active_2022, ).
@@
Figure 1. Basic PC framework, in which the error (i.e., the discrepancy) between the actual input from lower layers and the predicted value is the information propagated upwards through the network. Figure adapted from (rao_predictive_1999, ).
@@
Although PC was partly inspired by ML methods (dayan_helmholtz_1995, ; rao_predictive_1999, ), within the ML community it appears to have gone mostly unnoticed until recently. In 2017, Whittington & Bogacz applied PCNs to supervised learning tasks, revealing notable properties that connect them to traditional feedforward neural networks (FNNs). First, PCNs become equivalent to FNNs during inference (testing) (whittington_approximation_2017, ; song_predictive_2021, ; frieder_non-convergence_2022, ). Second, PCN’s training algorithm, called inference learning (IL) can be related to backpropagation (BP), the workhorse training algorithm of modern deep learning. These properties sparked a surge of interest in PCNs (song_can_2020, ; salvatori_predictive_2021, ; salvatori_reverse_2022, ; millidge_predictive_2022, ; millidge_theoretical_2022, ), suggesting IL as a more biologically plausible alternative for training deep learning architectures.
@@
Until recently, implementations of PCNs were more computationally intensive than equivalent BP-trained networks, helping to explain the lack of wider adoption of IL in ML applications. However, recent work has achieved marked improvements in performance, showing it can achieve higher efficiency than BP with sufficient parallelization (salvatori_stable_2023, ; alonso_understanding_2024, ). This is because IL’s computations use only locally available information, such that the serial updates inherent to BP can be avoided. If sufficiently parallelized, this means that computation time no longer scales with depth in PCNs (salvatori_stable_2023, ). As FNNs become increasingly large (deep), this could provide a marked advantage compared to BP, and further suggests PCNs are a promising candidate for use on neuromorphic hardware (millidge_predictive_2022-1, ; salvatori_stable_2023, ).
@@
While recent work has compared PCNs to FNNs, in particular focusing on supervised learning, fundamentally PCN is a probabilistic (Bayesian) model, naturally formulated for unsupervised learning. In fact, even before it was used within ML, the PCN of Rao & Ballard was phrased in terms familiar to the ML community (rao_predictive_1999, ; marino_predictive_2022, ). Specifically, it is defined by a graphical model with latent (unobserved) variables, trained using generalized Expectation Maximization (friston_learning_2003, ; bishop_pattern_2006, ), a general procedure for maximum likelihood in latent variable models (also used, e.g., in $k$-means clustering) (dempster_maximum_1977, ; neal_view_1998, ; bishop_pattern_2006, ). Its objective (i.e., cost function) is the complete data log-likelihood or variational free energy when seen as a variational inference procedure (friston_theory_2005, ; friston_free-energy_2010, ; bishop_pattern_2006, ). Hence, from this perspective, PCNs are most appropriately compared not to FNNs, but to techniques associated with generative modeling (e.g., variational autoencoders, diffusion models (prince_understanding_2023, )) and classic latent variable models (e.g., probabilistic PCA (bishop_pattern_2006, ), factor analysis (goodfellow_deep_2016, )). In other words, PC can be seen both as a learning algorithm contrasted with BP, and as a probabilistic latent variable model comparable to generative models.
In PCNs, the difference between supervised and unsupervised learning comes down to the direction of predictions within the network. In the supervised setting, these flow from data to the labels, while in the unsupervised setting, they flow towards the data (fig. 1). Recent work (salvatori_learning_2022, ) showed how this direction can be extended to a broader notion of topology by defining IL on any graph, known as PC graphs. This allows the training of non-hierarchical, brain-like structures, going beyond hierarchical to heterarchical prediction. This means PCNs effectively encompass FNNs, allowing the definition of a more general class of neural networks. While we are unaware of a formal proof in the literature, one expects that these networks share with FNNs the appealing property of being universal function approximators, but they also allow a new collection of structures untrainable with BP to be studied. Although the study of such structures is in its infancy, this means they form a superset of traditional ANNs (fig. 2). This understanding of PC graphs as generalized ANNs follows from (salvatori_learning_2022, ), but has not been previously phrased in this way.
@@
The different perspectives on PCNs for ML (fig. 3) are warranted by their mathematical structure, as outlined in (friston_hierarchical_2008, ; friston_free-energy_2010, ; buckley_free_2017, ; marino_predictive_2022, ). Substantial progress has been made since these works, however, both theoretically and in applying PCNs to larger datasets. Although some of this progress was surveyed in (millidge_predictive_2022, ; millidge_predictive_2022-1, ), and more recently in (salvatori_brain-inspired_2023, ; taniguchi_world_2023, ), these works favor broad coverage over mathematical detail. Additionally, the only tutorial on PCNs dates from 2017 (bogacz_tutorial_2017, ), before most work on PC in ML was published. Hence, there is a need for a more comprehensive formal specification, and a tutorial aimed at ML practitioners. Our work aims to provide both in the form of a combined tutorial/survey paper, as well as to clarify some minor mathematical errors in the literature.
@@
We focus on PC as a neuro-inspired branch of ML, meaning we choose not to discuss the large body of work on PC in neuroscience, psychology, philosophy, active inference, and robotics (see aforementioned references and (lanillos_active_2021, ; parr_active_2022, ) for reviews). Specifically, we do not discuss the neurologically-relevant issue of biological plausibility, often discussed in the PC literature (see e.g. (golkar_constrained_2022, ; song_inferring_2024, )).
@@
  • Summarize and provide structure for recent theoretical and empirical results on PCNs, integrating different perspectives (fig. 3), as both a unified basis for future work, and an accessible starting-point for machine learning practitioners.
  • Explain connections with existing methods in ML, some of which have remained unexposed in recent literature. In particular, we highlight how the structure of PCNs and PC graphs forms a mathematical superset of ANNs. This follows from earlier work (salvatori_learning_2022, ), but to our knowledge was not yet pointed out as a general conception of PCNs.
Our aim is to provide a technical reference while remaining accessible, providing a useful starting point for researchers seeking to understand PCNs and pursue future work in this area.
@@
The three perspectives in fig. 3 define the structure of this work. While interconnected, making a distinction facilitates an understanding of PCNs. Section 2 interprets PCNs as generalized ANNs (fig. 2),which should provide a familiar context for many ML researchers. Section 3 discusses PC as probabilistic latent variable models, and section 4 as a learning algorithm, focusing on theoretical results on IL as compared to BP. The literature on PCNs is organized in Table 1 by network type (section 2), which we survey throughout this work. We conclude in section 5, with appendices including an extended discussion of section 3 (appendix A), complementary proofs (appendix B.2), and a discussion of computational complexity (appendix C).
@@
Background Existing reviews (bogacz_tutorial_2017, ; millidge_predictive_2022, ; millidge_predictive_2022-1, ; marino_predictive_2022, ; salvatori_brain-inspired_2023, ) 1
PhD theses (millidge_applications_2021, ; song_predictive_2021, ; <salvatori_thesis>, ; alonso_energy-based_2023, )
Supervised learning IL’s relation to BP (whittington_approximation_2017, ; song_can_2020, ; salvatori_predictive_2021, ; millidge_predictive_2022, ; rosenbaum_relationship_2022, ; salvatori_reverse_2022, ; zahid_curvature-sensitive_2023, ; millidge_backpropagation_2022, ) 2, 4
IL in its natural regime (alonso_theoretical_2022, ; millidge_theoretical_2022, ; innocenti_understanding_2023, ; song_inferring_2024, ; frieder_non-convergence_2022, ; millidge_backpropagation_2022, ; frieder_bad_2024, ; salvatori_stable_2023, ; alonso_understanding_2024, )
Extensions (salvatori_learning_2022, ; salvatori_stable_2023, ; byiringiro_robust_2022, ; pinchetti_predictive_2022, )
Supervised generative modeling (kinghorn_preventing_2023, ; salvatori_learning_2022, )
Unsupervised learning Unsupervised generative modeling (rao_predictive_1999, ; salvatori_associative_2021, ; ororbia_neural_2022, ; ofner_generalized_2022, ; zahid_predictive_2022, ; zahid_sample_2024, ) 2.5, 3
Other variations and applications Memory models (tang_recurrent_2023, ; tang_sequential_2023, ; li_modeling_2023, )
Temporal prediction (jiang_dynamic_2024, ; millidge_predictive_2024, )
Other (boutin_iterative_2020, ; tscshantz_hybrid_2023, ; salvatori_causal_2023, )
Other uses of PC in ML Neural generative coding (NGC) (ororbia_neural_2022, ; ororbia_lifelong_2022, ; ororbia_backprop-free_2022, ; ororbia_convolutional_2023, ; ororbia_spiking_2023, )
PC-Inspired ANNs (chalasani_deep_2013, ; lotter_deep_2017, ; han_deep_2018, ; ye_anopcn_2019, ; rathjens_classification_2024, ; choksi_brain-inspired_2020, ; choksi_predify_2021, ; pang_predictive_2021, ; alamia_role_2021, ; faye_mathematical_2023, )
@@ 2. PCNs as Generalized ANNs
The theory of PC as a machine learning algorithm is often derived from variational inference in a hierarchical Gaussian latent variable model. The resulting equations bear a lot of similarity to those of artificial neural networks (ANNs). In particular, PCNs of a particular type, sometimes called discriminative PCNs, are in an important sense equivalent to feedforward neural network during inference (testing). Since we believe ANNs to be more familiar context for most ML researchers and practitioners than variational inference, we will here take ANNs as a starting point, with the aim of making the generalization to PCNs more transparent.
Thus, we review briefly the theory of the simplest ANN: a feedforward neural network (FNN; also known as a multilayer perceptron) in a supervised learning context, as taught in introductory ML classes, after which we review discriminative PCNs. To understand similarities and differences between both networks, we discuss in detail their training procedure, learning algorithm, testing procedure, and structure.
Problem Setup. We are given a dataset of $N$ labelled samples $\{\bm{x}^{(n)},\bm{y}^{(n)}\}_{n=1}^{N}$ split into a training set and a test set, where $\bm{x}^{(n)}\in X$ is a datapoint and $\bm{y}^{(n)}\in Y$ its corresponding label (bishop_pattern_2006, ), defined on input and output domains $X$, $Y$, respectively.

### v2

# skipabove=10pt,skipbelow=10pt

@@ Abstract.
Recent years have witnessed a growing call for renewed emphasis on neuroscience-inspired approaches in artificial intelligence research, under the banner of NeuroAI. A prime example of this is predictive coding networks (PCNs), based on the neuroscientific framework of predictive coding. This framework views the brain as a hierarchical Bayesian inference model that minimizes prediction errors through feedback connections. Unlike traditional neural networks trained with backpropagation (BP), PCNs utilize inference learning (IL), a more biologically plausible algorithm that explains patterns of neural activity that BP cannot. Historically, IL has been more computationally intensive, but recent advancements have demonstrated that it can achieve higher efficiency than BP with sufficient parallelization. Furthermore, PCNs can be mathematically considered a superset of traditional feedforward neural networks (FNNs), significantly extending the range of trainable architectures. As inherently probabilistic (graphical) latent variable models, PCNs provide a versatile framework for both supervised learning and unsupervised (generative) modeling that goes beyond traditional artificial neural networks. This work provides a comprehensive review and detailed formal specification of PCNs, particularly situating them within the context of modern ML methods. Additionally, we introduce a Python library (PRECO) for practical implementation. This positions PC as a promising framework for future ML innovations.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Neuroscience-inspired approaches to machine learning (ML) have a long history within artificial intelligence research (Rosenblatt, 1958; Hassabis et al., 2017; Macpherson et al., 2021). Despite the remarkable empirical advances in ML capabilities in recent years, biological learning is still superior in many ways, such as flexibility and energy efficiency (Macpherson et al., 2021). As such, recent years have seen a growing call for renewed emphasis on neuroscience-inspired approaches in AI research, known as NeuroAI (Zador et al., 2023; noa, 2024). This is exemplified by the rising popularity of the predictive coding (PC, also known as predictive processing) framework in computational neuroscience (Rao and Ballard, 1999; Friston, 2010; Ficco et al., 2021) and its recent entrance into the field of machine learning (Whittington and Bogacz, 2017; Millidge et al., 2022b). PC represents a Bayesian perspective on how the brain processes information, emphasizing the role of probabilistic models and minimization of prediction errors in perception and learning (Rao and Ballard, 1999; Friston, 2008, 2010; Ficco et al., 2021). At its core, it suggests that the brain continually generates predictions about sensory input and updates these predictions based on incoming sensory data. By removing the predictable components, this reduces redundancy in the information processing pipeline (Huang and Rao, 2011).
@@
Central to PC is the concept of hierarchical prediction: the lowest level of the hierarchy represents sensory data, and each higher level attempts to predict neural activity in the layer below (Rao and Ballard, 1999). Prediction errors, arising from discrepancies between actual and predicted activity, are propagated upward through the hierarchy, while the predictions from higher levels are propagated downwards via feedback connections; see fig. 1. In the influential work by Rao & Ballard (1999), the authors showed that such a predictive coding network (PCN) can learn statistical regularities in the input and explain several neural responses in the visual cortex (Shain et al., 2020; Okada et al., 2018; Ficco et al., 2021; Wang et al., 2023; Caucheteux et al., 2023). Since then, the PC framework has found wide adoption across neuroscience, psychology, and philosophy (Schrimpf et al., 2021; Clark, 2016). It is often seen as part of ‘active inference’, a broader umbrella term which aims to provide an integrated perspective on brain, cognition, and behavior (Parr et al., 2022).
@@
Figure 1. Basic PC framework, in which the error (i.e., the discrepancy) between the actual input from lower layers and the predicted value is the information propagated upwards through the network. Figure adapted from (Rao and Ballard, 1999).
@@
Although PC was partly inspired by ML methods (Dayan et al., 1995; Rao and Ballard, 1999), within the ML community it appears to have gone mostly unnoticed until recently. In 2017, Whittington & Bogacz applied PCNs to supervised learning tasks, revealing notable properties that connect them to traditional feedforward neural networks (FNNs). First, PCNs become equivalent to FNNs during inference (testing) (Whittington and Bogacz, 2017; Song, 2021; Frieder and Lukasiewicz, 2022). Second, PCN’s training algorithm, called inference learning (IL) can be related to backpropagation (BP), the workhorse training algorithm of modern deep learning. These properties sparked a surge of interest in PCNs, suggesting IL as a more biologically plausible alternative for training deep learning architectures (Song et al., 2020; Salvatori et al., 2021b, 2022b; Millidge et al., 2022a, 2023b).
Recent work (Song et al., 2024) showed that IL differs from BP by using prospective configuration. According to this principle, neural activities across the network change before synaptic weights are modified, in order for neurons to better predict future inputs – making activity changes ‘prospective’. (Song et al., 2024) showed that this principle naturally explains several patterns of neural activity and behavior in human and animal learning experiments which are unexplainable by BP – providing additional evidence for its biological plausibility. This property has been linked to IL’s sensitivity to second-order information in the loss landscape (Innocenti et al., 2023; Alonso et al., 2024), which leads to faster convergence and enhanced performance on certain learning tasks, like continual learning and online learning (Song et al., 2024; Alonso et al., 2022; Ororbia and Kifer, 2022; Innocenti et al., 2023).
@@
Until recently, implementations of PCNs were more computationally intensive than equivalent BP-trained networks, helping to explain the lack of wider adoption of IL in ML applications. However, recent work has achieved marked improvements in performance, showing it can achieve higher efficiency than BP with sufficient parallelization (Salvatori et al., 2024b; Alonso et al., 2024). This is because IL’s computations use only locally available information, such that the serial updates inherent to BP can be avoided. If sufficiently parallelized, this means that computation time no longer scales with depth in PCNs (Salvatori et al., 2024b). As FNNs become increasingly large (deep), this could provide a marked advantage compared to BP, and further suggests PCNs are a promising candidate for use on neuromorphic hardware (Millidge et al., 2022b; Salvatori et al., 2024b).
@@
While recent work has compared PCNs to FNNs, in particular focusing on supervised learning, fundamentally the PCN is a probabilistic (Bayesian) model, naturally formulated for unsupervised learning. In fact, even before it was used within ML, the PCN of Rao & Ballard was phrased in terms familiar to the ML community (Rao and Ballard, 1999; Marino, 2022). Specifically, it is defined by a graphical model with latent (unobserved) variables, trained using generalized Expectation Maximization (Friston, 2003; Bishop, 2006), a general procedure for maximum likelihood in latent variable models (also used, e.g., in $k$-means clustering) (Dempster et al., 1977; Neal and Hinton, 1998; Bishop, 2006). Its objective (i.e., cost function) is the complete data log-likelihood or variational free energy when seen as a variational inference procedure (Friston, 2005, 2010; Bishop, 2006). Hence, from this perspective, PCNs are most appropriately compared not to FNNs, but to techniques associated with generative modeling (e.g., variational autoencoders, diffusion models (Prince, 2023)) and classic latent variable models (e.g., probabilistic PCA (Bishop, 2006), factor analysis (Goodfellow et al., 2016)). In other words, PC can be seen both as a learning algorithm contrasted with BP, and as a probabilistic latent variable model comparable to generative models.
@@
In PCNs, the difference between supervised and unsupervised learning comes down to the direction of predictions within the network. In the supervised setting, these flow from data to the labels, while in the unsupervised setting, they flow towards the data (fig. 1). Recent work (Salvatori et al., 2022a) showed how this direction can be extended to a broader notion of topology by defining IL on any graph, known as PC graphs. This allows the training of non-hierarchical, brain-like structures, going beyond hierarchical to heterarchical prediction. This means PCNs effectively encompass FNNs, allowing the definition of a more general class of neural networks. While we are unaware of a formal proof in the literature, one hence expects that these networks share with FNNs the appealing property of being universal function approximators. Simultaneously, they allow a new collection of structures untrainable with BP to be studied. Although the study of such structures is in its infancy, this means they form a superset of traditional ANNs (fig. 2). This understanding of PC graphs as generalized ANNs follows from (Salvatori et al., 2022a), but has not been previously phrased in this way.
The different perspectives on PCNs for ML (fig. 3) are warranted by their mathematical structure, as outlined in (Friston, 2008, 2010; Buckley et al., 2017; Marino, 2022). Substantial progress has been made since these works, however, both theoretically and in applying PCNs to larger datasets. Although some of this progress was surveyed in (Millidge et al., 2022a, b), and more recently in (Salvatori et al., 2023; Taniguchi et al., 2023), these works favor broad coverage over mathematical detail. Additionally, the only tutorial on PCNs dates from 2017 (Bogacz, 2017), before most work on PC in ML was published. Hence, there is a need for a more comprehensive formal specification, and a tutorial aimed at ML practitioners. Our work aims to provide both in the form of a combined tutorial/survey paper, as well as to clarify some minor mathematical errors in the literature.
@@
We focus on PC as a neuro-inspired branch of ML, meaning we choose not to discuss the large body of work on PC in neuroscience, psychology, philosophy, active inference, and robotics (see aforementioned references and (Lanillos et al., 2021; Parr et al., 2022) for reviews). Specifically, we do not discuss the neurologically-relevant issue of biological plausibility, often discussed in the PC literature (see e.g. (Golkar et al., 2022; Song et al., 2024)).
@@
  • Summarize and provide structure for recent theoretical and empirical results on PCNs, integrating different perspectives (fig. 3), as both a unified basis for future work, and starting-point for machine learning practitioners. Our aim is to provide a technical reference while remaining accessible.
  • Explain connections with existing methods in ML, some of which have remained unexposed in recent literature. In particular, we highlight how the structure of PCNs and PC graphs forms a mathematical superset of ANNs. This follows from earlier work (Salvatori et al., 2022a), but to our knowledge was not yet pointed out as a general conception of PCNs.
  • Providing a Python library (PRECO [^1] Available at https://github.com/bjornvz/PRECO.) using PyTorch that implements PCNs and PC graphs, as a hands-on tutorial.
@@
The three perspectives in fig. 3 define the structure of this work. While interconnected, making a distinction facilitates an understanding of PCNs. Section 2 interprets PCNs as generalized ANNs (fig. 2), which should provide a familiar context for many ML researchers. Section 3 discusses PCNs as probabilistic latent variable models, and section 4 discusses work focusing on IL, i.e. PC as a learning algorithm. The literature on PCNs is organized in Table 1 by network type (section 2), which we survey throughout this work. We conclude in section 5, with appendices including an extended discussion of section 3 (appendix A), complementary proofs (appendix B.2), and a discussion of computational complexity (appendix C).
@@
Background Existing reviews (Bogacz, 2017; Millidge et al., 2022a, b; Marino, 2022; Salvatori et al., 2023) 1
PhD theses (Millidge, 2021; Song, 2021; Salvatori, 2022; Alonso, 2023)
Discriminative PCNs IL’s relation to BP (Whittington and Bogacz, 2017; Song et al., 2020; Salvatori et al., 2021b; Millidge et al., 2022a; Rosenbaum, 2022; Salvatori et al., 2022b; Zahid et al., 2023b; Millidge et al., 2023a) 2, 4
IL in its natural regime (Salvatori et al., 2023; Millidge et al., 2023b; Innocenti et al., 2023; Song et al., 2024; Frieder and Lukasiewicz, 2022; Millidge et al., 2023a; Frieder et al., 2024; Salvatori et al., 2024b; Alonso et al., 2024)
Extensions (Salvatori et al., 2022a, 2024b; Byiringiro et al., 2022; Pinchetti et al., 2022)
Generative PCNs Supervised learning (Kinghorn et al., 2023; Salvatori et al., 2022a) 2.5, 3
Unsupervised learning (Rao and Ballard, 1999; Salvatori et al., 2021a; Ororbia and Kifer, 2022; Ofner et al., 2022; Zahid et al., 2023a, 2024)
Other variations and applications Memory models (Tang et al., 2023b, a; Li et al., 2023)
Temporal prediction (Jiang and Rao, 2024; Millidge et al., 2024)
Other (Boutin et al., 2020; Tscshantz et al., 2023; Salvatori et al., 2024a)
Other uses of PC in ML Neural generative coding (NGC) (Ororbia and Kifer, 2022; Ororbia et al., 2022; Ororbia and Mali, 2022, 2023; Ororbia, 2023)
PC-Inspired ANNs (Chalasani and Principe, 2013; Lotter et al., 2017; Han et al., 2018; Ye et al., 2019; Rathjens and Wiskott, 2024; Choksi et al., 2020, 2021; Pang et al., 2021; Alamia et al., 2023; Faye et al., 2023)
@@ 2. PCNs as Generalized ANNs
The theory of PC as a machine learning algorithm is often derived from variational inference in a hierarchical Gaussian latent variable model. The resulting equations bear a lot of similarity to those of artificial neural networks (ANNs), however. In particular, PCNs of a particular type, sometimes called discriminative PCNs, are in an important sense equivalent to feedforward neural network during inference (testing). Since we believe ANNs to be more familiar context for most ML researchers and practitioners than variational inference, we will here take ANNs as a starting point, with the aim of making the generalization to PCNs more transparent.
Thus, we review briefly the theory of the simplest ANN: a feedforward neural network (FNN; also known as a multilayer perceptron) in a supervised learning context, as taught in introductory ML classes, after which we review discriminative PCNs. To understand similarities and differences between both networks, we discuss in detail their training procedure, learning algorithm, structure, and testing procedure. [^2] At this point, we note briefly that a second usage of term ‘PCN’ exists, referring to backprop-trained ANNs with an architecture that takes inspiration from PC, e.g. includes separate top-down and bottom-up FNNs (cf. ‘PC-inspired ANNs’ in table 1).
Problem Setup.
We are given a dataset of $N$ labelled samples $\{\bm{x}^{(n)},\bm{y}^{(n)}\}_{n=1}^{N}$ split into a training set and a test set, where $\bm{x}^{(n)}\in X$ is a datapoint and $\bm{y}^{(n)}\in Y$ its corresponding label, defined on input and output domains $X$, $Y$, respectively (Bishop, 2006).